What is the final temperature of a sample of a nickel if 328 j of energy is added to a 16.7g sample at an initial temperature of 24.4c
1
Expert's answer
2019-05-24T02:00:27-0400

C(Ni) = 500 J/(kg*K)

m(Ni) = 0.0167 kg

Q = 328 J

T = T0+ (Q/(Cm)) = 24.4 +39.29 = 63.68 C